An electronics manufacturing process has a scheduled mean completion time of 70 minutes. It is claimed that under new management, the mean completion time,m , is less than 70 minutes. To test this claim, a random sample of 15 completion times under new management was taken.
The sample had a mean completion time of 68 minutes and a standard deviation of 10 minutes. Assume that the population of completion times under new management is normally distributed. At the 0.05 level of significance, can it be concluded that the mean completion time,m , under new management is less than the scheduled mean?
Perform a one-tailed test. Then fill in the table below.

Answer by stanbon(75887)

An electronics manufacturing process has a scheduled mean completion time of 70 minutes. It is claimed that under new management, the mean completion time,m , is less than 70 minutes. To test this claim, a random sample of 15 completion times under new management was taken.
The sample had a mean completion time of 68 minutes and a standard deviation of 10 minutes. Assume that the population of completion times under new management is normally distributed. At the 0.05 level of significance, can it be concluded that the mean completion time,m , under new management is less than the scheduled mean?
Perform a one-tailed test.
--------
Ho: u >= 70
Ha: u < 70 (claim)
---------------------
t(68) = (68-70)/[10/sqrt(15)] = -0.7750
---
p-value = P(t < -0.7750 when df=14) = 0.2256
===========================
Conclusion: Since the p-value is greater than 5%,
fail to reject Ho.
The test results do not support the claim.
============================
Cheers,
Stan H.